# Effective Two-Stage Knowledge Transfer for Multi-Entity Cross-Domain Recommendation

### v1

@@ 3.2.3. Common Knowledge Extractor (CKE)
After acquiring the aligned entity vectors from HFA, we need to extract the shared knowledge from products and content posts. In theory, any network architecture with shared-private parameters can serve as the common knowledge extractor for MKT, such as STAR (Sheng et al., 2021), MMoE (Ma et al., 2018), or PLE (Tang et al., 2020). Here, we opt for the PLE model structure as our common knowledge extractor, referred to as CKE.
More specifically, we use $f_{CKE}$ to represent the network operations of CKE. As a result, the output vectors $\mathbf{q}^{s}$ and $\mathbf{q}^{t}$ from HFA will then transform into high-order knowledge extracted from user-entity interactions, as shown in the following formula:
@@
3.2.4. Polarized Distribution Loss
During joint training, it is hard to measure whether these shared parameters indeed learn the common knowledge, and to what extent they are doing so. To address this issue, we propose an auxiliary task that enforces the Common Knowledge Extractor (CKE) to surely learn the common knowledge. We name the corresponding loss function as Polarized Distribution Loss (PDL), which will be discussed lately in the below paragraphs.
Note that along with the CKE, we have also designed independent knowledge extractors for different entities: the Source Knowledge Extractor (SKE) and the Target Knowledge Extractor (TKE). These entity-independent knowledge extractors can be as simple as an MLP model. For simplicity, we use $f_{SKE}$ and $f_{TKE}$ to denote the knowledge extractors for the source and target entity, respectively. Then, we can obtain the independent high-order knowledge $\mathbf{g}^{s}_{ind}$ for the source entity, and the independent high-order knowledge $\mathbf{g}^{t}_{ind}$ for the target entity through the following formula:
@@ 3.2.4. Polarized Distribution Loss
We set the output vector dimension of SKE and TKE to be the same as that of CKE. The reason is that once we obtain independent and common knowledge, we want them to be as different as possible. If the corresponding knowledge extractors capture similar knowledge, the information gain for the target domain would be minimal. To this end, we adopt the polarized distribution loss to minimize the similarity between the independent knowledge and the common knowledge. The cosine function is a commonly adopted similarity function for this purpose. This process can be illustrated by the below formula:
@@
$\|\cdot\|_{2}$ means the length of a vector in the Euclidean space. With the above notations, the polarized distribution loss is as follows:
(12) $\begin{split}Loss_{pdl}&=Loss_{src\_sim}+Loss_{tgt\_sim}\\ &=cosine(\mathbb{g}^{s}_{com},\mathbb{g}^{s}_{ind})\\ &+cosine(\mathbb{g}^{t}_{com},\mathbb{g}^{t}_{ind})\end{split}$
@@
(13) $\begin{split}Loss_{MEM}&=Loss_{src}+Loss_{tgt}+\gamma Loss_{pdl}\\ &=\sum_{x_{i}\in X^{s}}^{N_{s}}L(y_{i}^{s},f_{MEM}(x_{i},\Theta_{MEM}))\\ &+\sum_{x_{i}\in X^{t}}^{N_{t}}L(y_{i}^{t},f_{MEM}(x_{i},\Theta_{MEM}))\\ &+\gamma Loss_{pdl}\end{split}$
where $x_{i}$ stands for a data sample either from the source or the target domain, $\gamma$ is the hyper-parameter (we set the $\gamma$ parameter to 0.1 throughout the whole paper) for $Loss_{pdl}$, $f_{MEM}$ represents the MEM, $L$ denotes the loss function, and $\Theta_{MEM}$ indicate the trainable parameters in MEM.
@@ 3.3. Target-Entity Model Fine-tuning
In the first stage, we train the MEM model with mixed data samples from both the source entity and the target entity. However, we notice that the number of data samples for the source entity is three times more than that of the target entity. This may result in the MEM being dominated by the source entity and may be sub-optimal if it is used for online serving. To overcome this issue, we introduce a second stage fine-tuning process: with the obtained MEM model from the first stage, we integrate the corresponding knowledge into a Target-Entity Model(TEM) for fine-tuning. By doing so, we may achieve better model performance.
To be more specific, we freeze the parameters of MEM during the transfer of knowledge from MEM to TEM. Only target entity samples are required for TEM fine-tuning. To ensure that TEM keeps the entity-specific ability, we only transfer knowledge vectors with the same semantics. These include the user profile feature embedding vector, user behavior sequence vector, entity-shared feature embedding vector, and the user-entity interaction knowledge vector extracted by the CKE module.
During the second-stage fine-tuning, we only feed the target entity data into both MEM and TEM. Here, we utilize the vectors $\mathbb{v}_{u}^{t}$, $\mathbb{v}_{seq}^{t}$, $\mathbb{v}_{shared}^{t}$, $\mathbb{g}_{com}^{t}$ from MEM, and the user feature vector $\mathbb{v}_{u}^{TEM}$, user behavior sequence vector $\mathbb{v}_{seq}^{TEM}$, entity-shared feature vector $\mathbb{v}_{{}_{s}hared}^{TEM}$, and user-entity cross vector $\mathbb{g}_{cross}^{TEM}$ from TEM.
Considering that the common knowledge in the MEM is not entirely applicable to TEM, it is necessary to filter out irrelevant knowledge. This helps to ensure that the resulting common knowledge is truly useful for the target domain. To achieve this, we use an adaptive gating structure (such as GLU (Dauphin et al., 2017)) when plugging the knowledge vector from MEM. This gating structure helps to filter out irrelevant knowledge and allows us to combine the filtered common knowledge with the vectors from TEM. For example, the user-entity interaction vector has a gate denoted as $f_{cross\_gate}$, and the formula for this gate is as follows.
(14) $\mathbb{g}_{com\_gate}^{t}=f_{cross\_gate}(\mathbb{g}_{com}^{t})$
(15) $\mathbb{g}_{cross\_final}^{TEM}=\mathbb{g}_{cross}^{TEM}+\mathbb{g}_{com\_gate}^{t}$
After receiving knowledge from the multi-entity model, we get the final user-entity interaction vector for TEM. All of the knowledge vectors such as $\mathbb{v}_{u\_final}^{TEM}$, $\mathbb{v}_{seq\_final}^{TEM}$ and $\mathbb{v}_{shared\_final}^{TEM}$ can be obtained in the same manner.
The overall loss of TEM can be formulated as follows:
@@
where $x_{i}\in X^{t}$ denotes a data sample from the target domain, $f_{TEM}$ represents TEM, $L$ denotes the cross-entropy loss adopted in this paper, and $\Theta_{TEM}$ is the trainable parameter for TEM.
@@ 4. Experiment
In this section, we conduct extensive experiments to understand the effectiveness of the MKT algorithm.
@@ 4.1.1. Datasets
Due to the lack of a publicly available large-scale multi-entity dataset from the industrial recommender systems, we construct experimental data with the impression log from our production system. We collect 31 days of user interaction data from the Xianyu homepage recommendation system. As mentioned before, 90% of them are related to selling products and the remaining 10% is regarding the content posts. Since the magnitude of user interaction data is beyond processing, we keep all positive samples (i.e., user clicks) while conducting sampling for negative ones (the sampling ratios are different for product and post). In total, we obtain 9.3 billion data samples for the source entity and 3.1 billion data samples for the target entity. The overall statistics are summarized in Table 1. In the below experiments, we use 30 days of data for training and the remaining one day of data for testing.
Table 1. Overall statistics of the collected production dataset.
Source entity (products) Target entity (posts)
Users 130M 130M
Items 290M 4M
Data samples 9.3B 3.1B
@@ 4.1.2. Baselines
We include both single-domain and cross-domain recommendation algorithms for comparison. The cross-domain approaches are further divided into two categories: the joint training method (i.e., the below MMOE, PLE, STAR, CoNet, and MiNet algorithms), and the pre-training & fine-tuning method (i.e., the below Finetune and CTNet models).
@@ 4.1.3. Implementation Details
Embedding dimensions are set to 8 for all of the sparse features. For the above-mentioned single-domain models and the Target Entity Model (see Section 3.3), the number of layers for the Deep Neural Network is set to 4, with the number of hidden units in each layer set to [1024, 512, 64, 1]. As for the joint training approaches and Multi-Entity Model (see Section 3.2), we employ a two-layer shared network, with a configuration of hidden units as [1024, 512]. The shared layers are then followed by domain-specific towers for each domain, with each tower consisting of two layers with hidden units [64, 1]. For MMOE, we set the number of expert networks to 3. Since CTNet can not deal with multiple entities, we have to discard the entity-specific features during pre-training.
@@ 4.2.1. Overall Evaluation
Table 2 provides the model performance for all of the baseline approaches and our proposed Multi-entity Knowledge Transfer (MKT) algorithm. We discover that single-domain recommendation algorithms are generally inferior to the cross-domain methods, indicating that the data from different domains indeed helps better model user interest. The joint-training methods outperform the single-domain models, but not by a large margin. Most of them only surpass FiBiNet by 0.2% in both GAUC and AUC metrics, and the GAUC of CoNet is even on par with FiBiNet. This, we hypothesize, is mainly ascribed to their inability to address the heterogeneous feature schemas across multiple entities, resulting in limited improvements.
Pre-training & fine-tuning approaches are in general better than the joint-training algorithms. This might come from two aspects: (1) because of the data distribution difference and feature misalignment, joint-training algorithms are difficult to resolve the negative knowledge transfer issue, and (2) with the second-stage fine-tuning on the target entity data, the pre-training & fine-tuning models can better adapt to the target entity.
As for the pre-training & fine-tuning methods, MKT achieves the best performance on both AUC and GAUC metrics. The most prominent difference is that MKT pre-trains a multi-entity base model with mixed source entity and target entity data, whereas the Finetune and CTNet models only utilize the data from the source entity. With the target entity data for pre-training, the model can capture the target entity distribution in ahead. Besides, through the well-designed Heterogeneous Feature Alignment (HFA) architecture and polarized distribution loss, MKT can effectively extract desired knowledge, and avoid local optima that might exist in Finetune and CTNet. Note that, here, MKT transfers user feature embedding, shared entity feature embedding, and the extracted common knowledge to the target entity model (see more details in Figure 2).
@@
Model AUC GAUC
MLP 0.7315 0.5683
Single-domain DIN 0.7335 0.5761
FiBiNet 0.7345 0.5824
MMoE 0.7367 0.5856
PLE 0.7363 0.5837
Joint-training STAR 0.7362 0.5845
CoNet 0.7357 0.5820
MiNet 0.7370 0.5863
Pre-training Finetune 0.7372 0.5897
& CTNet 0.7380 0.5872
Fine-tuning MKT 0.7398 0.5958
4.2.2. Model Performance on Different User Groups
To further understand the behavior of the proposed MKT model, we split users into three groups based on their activity levels (which are measured by the number of clicks over the past month). Particularly, users with fewer than 10 clicks are assigned to Group 1, users with [10, 30) clicks are put into Group 2, and the rest are in Group 3. Here, we select FiBiNet and Finetune for comparison. FiBiNet is the best-performing single-domain recommendation model. Finetune falls under the same pre-training & fine-tuning recommendation paradigm as MKT, and it is also our production baseline.
Table 3. Model performance on different user groups.
@@ 4.2.3. Online A/B Testing
We successfully deployed MKT in our production environment to serve the main traffic of Xianyu APP Homepage recommendation since October of 2023. In this section, we provide online results by comparing MKT with our production baseline. Our production baseline follows the pre-training & fine-tuning paradigm, in which a base model with only the source entity data is pre-trained, and then the resulting model is fine-tuned on the target entity data, i.e., the Finetune method in Section 4.1.2. We conduct an online A/B testing between MKT and our production baseline from October 1 to October 7, each with 5% randomly assigned traffic. Two key business metrics used for evaluation include the Click-Through Rate (CTR) and User Engagement Rate (UER)  [^3] UER is one of our top-line business metrics that measures whether a user has liked/commented/collected on a target item.. During the time of A/B testing, we have observed an increase of CTR metric by +4.13%, and UER Metric by +7.07%, which demonstrates the considerable business value of MKT.
@@ 4.3. Ablation Study
In this section, we conduct a set of ablation studies to further understand the importance of each component in MKT. Particularly, we examine three modules: the Target Entity Model fine-tuning (see Section 3.3), the Heterogeneous Feature Alignment (see Section 3.2.2), and the Polarized Distribution loss (see Section 3.2.4).
According to Table 4, removing any of the three modules can lead to degenerated model performances by a fair margin. By dropping the TEM fine-tune module, both AUC and GAUC metrics decrease dramatically, indicating the great importance of the second-stage model fine-tuning. MKT without the Polarize Distribution Loss module leads to a drop of 0.11% and 0.35% in the AUC and GAUC metrics, respectively. Similarly, after removing HFA, model performance drops by 0.16% in AUC and 0.88% in GAUC. The main reason is that HFA helps prescreen feature importance, which eventually ensures more effective knowledge extraction in later modules.
Table 4. Model performance on different ablation studies.
Ablations AUC GAUC
value Imporv. value Improv.
w/o TEM Finetune 0.7374 -0.33% 0.5884 -1.26%
w/o PDL 0.7390 -0.11% 0.5937 -0.35%
w/o HFA 0.7386 -0.16% 0.5906 -0.88%
Full MKT 0.7398 - 0.5958 -

### v2

@@ 3.2.3. Common Knowledge Extractor (CKE)
After acquiring the aligned entity vectors from HFA, we need to extract the shared knowledge from source domain and target domain. In theory, any network architecture with shared parameters can serve as the common knowledge extractor for MKT, such as MLP or MMoE (Ma et al., 2018). Here, we opt for the MMoE model structure as our common knowledge extractor, referred to as CKE.
More specifically, we use $f_{CKE}$ to represent the network operations of CKE. Then, we input the output vectors $\mathbf{q}^{s}$ and $\mathbf{q}^{t}$ into CKE, as shown in the following formula:
@@
3.2.4. Polarized Distribution Loss (PDL)
During joint training, it is hard to measure whether these shared parameters indeed learn the common knowledge, and to what extent they are doing so. To address this issue, we propose an auxiliary task that enforces the Common Knowledge Extractor (CKE) to surely learn the common knowledge. We name the corresponding loss function as Polarized Distribution Loss. Note that along with the CKE, we have also designed independent knowledge extractors for different entities: the Source Knowledge Extractor (SKE) and the Target Knowledge Extractor (TKE). These entity-independent knowledge extractors can be as simple as MLP. For simplicity, we use $f_{SKE}$ and $f_{TKE}$ to denote the knowledge extractors for the source and target entity respectively. Then, we can obtain the independent knowledge $\mathbf{g}^{s}_{ind}$ and $\mathbf{g}^{t}_{ind}$ for the source and target entity through the following formula:
@@ 3.2.4. Polarized Distribution Loss (PDL)
We set the output vector dimension of SKE and TKE to be the same as that of CKE. We want the independent and common knowledge to be as different as possible. If the corresponding knowledge extractors capture similar knowledge, the information gain for the target domain would be minimal. So we adopt the PDL to minimize the similarity between them. The cosine function is a commonly adopted similarity function for this purpose. This process can be illustrated by the below formula:
@@
$\|\cdot\|_{2}$ means the length of a vector in the Euclidean space. With the above notations, the PDL is as follows:
(10) $\begin{split}Loss_{pdl}&=Loss_{src\_sim}+Loss_{tgt\_sim}\\ &=cosine(\mathbb{g}^{s}_{com},\mathbb{g}^{s}_{ind})+cosine(\mathbb{g}^{t}_{com},\mathbb{g}^{t}_{ind})\end{split}$
@@
(11) $\begin{split}Loss_{MEM}&=Loss_{src}+Loss_{tgt}+\gamma Loss_{pdl}\\ &=\sum_{x_{i}\in X^{s}}^{N_{s}}L(y_{i}^{s},f_{MEM}(x_{i},\Theta_{MEM}))\\ &+\sum_{x_{i}\in X^{t}}^{N_{t}}L(y_{i}^{t},f_{MEM}(x_{i},\Theta_{MEM}))+\gamma Loss_{pdl}\end{split}$
where $x_{i}$ stands for a data sample either from the source or the target domain, $\gamma$ is the hyper-parameter for $Loss_{pdl}$, $f_{MEM}$ represents the operators of MEM, $L$ denotes the loss function, and $\Theta_{MEM}$ indicate the trainable parameters in MEM.
@@ 3.3. Target-Entity Model Fine-tuning
In the first stage, we train the MEM with mixed data samples from both the source and target entity. However, we notice that the number of data samples for the source entity is far more than that of the target entity. This will result in the MEM being dominated by the source entity and encounters the negative transfer problem (Zhang and Yang, 2021). To overcome this issue, we introduce a second stage fine-tuning process: we connect the TEM with the obtained MEM model from the first stage, and integrate the corresponding knowledge of MEM into the TEM for fine-tuning. By doing so, we can achieve better model performance.
To be more specific, we freeze the parameters of MEM during the second stage. Only target entity samples are required for TEM fine-tuning. To ensure that TEM keeps the entity-specific ability, we only transfer knowledge vectors with the same semantics. These include the user profile feature embedding vector, user behavior sequence vector, entity-shared feature embedding vector, and the user-entity cross knowledge vector extracted by the CKE module.
Considering that the knowledge vectors in the MEM is not entirely applicable to TEM, it is necessary to filter out irrelevant knowledge. This helps to ensure that the resulting knowledge vectors are truly useful for the target domain. To achieve this, we use an adaptive gating structure (such as GLU (Dauphin et al., 2017)) when plugging the knowledge vectors from MEM. For example, the user-entity cross knowledge vector has a gate denoted as $f_{cross\_gate}$, after receiving knowledge from the MEM, we get the final user-entity cross vector for the TEM, and the formulas are as follows,
(12) $\mathbb{g}_{cross\_final}^{TEM}=\mathbb{g}_{cross}^{TEM}+f_{cross\_gate}(\mathbb{g}_{com}^{t})$
all of the knowledge vectors such as $\mathbb{v}_{u\_final}^{TEM}$, $\mathbb{v}_{seq\_final}^{TEM}$ and so on can be obtained in the same manner. The overall loss of TEM can be formulated as follows:
@@
where $x_{i}\in X^{t}$ denotes a data sample from the target domain, $f_{TEM}$ represents the operators of the TEM, $L$ denotes the cross-entropy loss adopted in this paper, and $\Theta_{TEM}$ is the trainable parameter for TEM.
@@ 4.1.1. Datasets
We conduct our experiments on public datasets and industrial datasets to compare our model and baseline models. The overall statistics are summarized in Table 1.
Tenrec dataset (Yuan et al., 2022): This dataset was released by Tencent and is the largest public dataset available for multi-entity cross-domain task that we could find. It is a user behavior log for two different feeds recommendation platforms of Tencent, namely, QQ BOW (QB) and QQ KAN (QK), including click, like, share and follow behaviors. An item in QK/QB can either be a news article or a video. Here, we use QK-video as the source entity and QK-article as the target entity because the number of data samples for QK is far more than that of QB and the same reason for video and article. Consider that QK-article does not contain negative samples for click behavior, we use like behavior type as positive feedback and no-like as negative feedback for QK-video and QK-article. As the dataset presents all interaction behaviors according to the time order, so we use the first 80% of data for training and the last 20% for testing.
Industrial dataset: We also construct experimental data with the impression log from our production system. We collect 31 days of user interaction data from the Xianyu homepage recommendation system, 75% of them are related to selling products and the remaining 25% is regarding the content posts. In total, we obtain 9.3 billion data samples for the source entity and 3.1 billion data samples for the target entity. In the below experiments, we use 30 days of data for training and the remaining one day of data for testing.
Table 1. Overall statistics of public and industrial datasets. Due to privacy concerns, we are unable to provide the positive sample ratio of the industrial dataset.
dataset Tenrec Industrial
domain video article product post
Users 5M 1.3M 130M 130M
Items 1.9M 0.2M 290M 4M
Data samples 142M 46M 9.3B 3.1B
positive sample ratio 0.036 0.018 \ \
@@ 4.1.2. Baselines
We include both single-domain and cross-domain recommendation algorithms for comparison. The cross-domain approaches are further divided into two categories: the joint training method and the pre-training & fine-tuning method (Finetune and CTNet).
@@ 4.1.3. Implementation Details
Embedding dimensions are set to 8 for all of the sparse features. For the above-mentioned single-domain models and the TEM (see Section 3.3), the number of layers for the Deep Neural Network is set to 4, with the number of hidden units in each layer set to [1024, 512, 64, 1] for the industrial dataset and [128, 64, 16, 1] for the public dataset. As for the joint training approaches and the MEM (see Section 3.2), we employ a two-layer shared network, with a configuration of hidden units as [1024, 512] for the industrial dataset and [128, 64] for the public dataset. The shared layers are then followed by domain-specific towers for each domain, with each tower consisting of two layers with hidden units [64, 1] for the industrial dataset and [16, 1] for the public dataset. For MMOE, we set the number of expert networks to 3. Since CTNet can not deal with multiple entities, we can only transfer the embedding of user and entity-shared features during fine-tuning. We set $\gamma$ to 0.01 for two datasets. The experiments are conducted multiple times and the mean value is taken as the model performance.
@@ 4.2.1. Overall Evaluation
Table 2 provides the model performance for all of the baseline models and our proposed Multi-entity Knowledge Transfer (MKT) model. We discover that single-domain recommendation algorithms are generally inferior to the cross-domain methods, indicating that the data from different domains indeed helps better model user interest.
Firstly, joint-training models generally outperform single-domain models. This superiority arises from joint-training’s ability to leverage common knowledge across multiple domains and enhance the model’s generalization capability. For instance, architectures like MMoE (Mixture-of-Experts) utilize shared expert networks and gating mechanisms to facilitate cross-domain knowledge sharing, resulting in improved recommendation performance across various domains. In contrast, single-domain models rely solely on data from a single domain, it is difficult to achieve good results in target domain with sparse data.
Secondly, the pre-training & fine-tuning methods further enhances performance beyond joint-training. This improvement is primarily attributed to its two-stage training strategy. During the pre-training stage, the model is trained on a great deal of source data, enabling it to learn universal knowledge representations. Subsequently, during the fine-tuning stage, the model is specifically optimized on target domain data, fine-tuning parameters to align with the target domain’s characteristics. This method not only preserves source domain knowledge but also effectively adapts to the target domain’s specific requirements, mitigating information conflicts and trade-offs that may arise in joint-training.
Finally, our proposed MKT model is further superior to the pre-training & fine-tuning methods on both AUC and GAUC metrics. This come from two aspects: (1) MKT performs pre-training on the mixed data of source domain and target domain, uses HFA to amplify the entity-spec features suitable for knowledge extracting, aligns different feature schemas, and uses PDL to extract cross-domain common knowledge more efficiently. (2) During the fine-tuning stage, MKT performs hierarchical transfer of knowledge with different semantics, including user features, entity-shared features, sequence features and user-entity cross vector, and uses gate to filter the knowledge from the pre-training stage to obtain more suitable parts.
@@
dataset Tenrec Industrial
Model AUC GAUC AUC GAUC
MLP 0.9337 0.6393 0.7315 0.5683
Single-domain DIN 0.9381 0.6441 0.7335 0.5761
FiBiNet 0.9374 0.6427 0.7345 0.5824
MMoE 0.9402 0.6461 0.7367 0.5856
PLE 0.9393 0.6473 0.7363 0.5837
Joint-training STAR 0.9382 0.6422 0.7362 0.5845
CoNet 0.9368 0.6413 0.7357 0.5820
MiNet 0.9420 0.6499 0.7370 0.5863
Pre-training Finetune 0.9456 0.6520 0.7372 0.5897
& CTNet 0.9439 0.6492 0.7380 0.5872
Fine-tuning MKT 0.9524 0.6583 0.7398 0.5958
Table 3. Model performance on different user groups of the industrial dataset.
@@
Table 4. Model performance on different ablation studies.
dataset Tenrec Industrial
Ablations AUC GAUC AUC GAUC
value Imporv. value Improv. value Imporv. value Improv.
w/o MEM Pretrain 0.9381 -1.50% 0.6441 -2.15% 0.7345 -0.71% 0.5824 -1.21%
w/o TEM Finetune 0.9447 -0.81% 0.6503 -1.21% 0.7374 -0.33% 0.5884 -1.26%
w/o PDL 0.9505 -0.19% 0.6569 -0.21% 0.7390 -0.11% 0.5937 -0.35%
w/o HFA 0.9477 -0.49% 0.6522 -0.92% 0.7386 -0.16% 0.5906 -0.88%
Full MKT 0.9524 - 0.6583 - 0.7398 - 0.5958 -
4.2.2. Model Performance on Different User Groups
To further understand the behavior of the proposed MKT model, we split users of the industrial dataset into three groups based on their activity levels (which are measured by the number of clicks over the past month). Here, we select FiBiNet and Finetune for comparison. FiBiNet is the best-performing single-domain recommendation model. Finetune falls under the same pre-training & fine-tuning recommendation paradigm as MKT, and it is also our production baseline.
@@ 4.2.3. Online A/B Testing
We successfully deployed MKT in our production environment to serve the main traffic of Xianyu APP Homepage recommendation since October of 2024. In this section, we provide online results by comparing MKT with our production baseline. Our production baseline follows the Finetune method in Section 4.1.2. We conduct an online A/B testing between MKT and our production baseline from October 1 to October 7, each with 5% randomly assigned traffic. Two key business metrics used for evaluation include the Click-Through Rate (CTR) and User Engagement Rate (UER)  [^2] UER is one of our top-line business metrics that measures whether a user has liked/commented/collected on a target item.. During the time of A/B testing, we have observed an increase of CTR metric by +4.13%, and UER Metric by +7.07%, which demonstrates the considerable business value of MKT.
@@ 4.3. Ablation Study
In this section, we conduct a set of ablation studies to further understand the importance of each component in MKT. Particularly, we examine four modules: the MEM pre-training (see Section 3.2), the TEM fine-tuning (see Section 3.3), the HFA (see Section 3.2.2), and the PDL (see Section 3.2.4).
According to Table 4, removing any of the four modules can lead to degenerated model performances by a fair margin. By dropping the MEM or TEM module, both AUC and GAUC metrics decrease dramatically, indicating the great importance of them. MKT without the PDL module leads to a drop of 0.11% and 0.35% in the AUC and GAUC metrics on industrial dataset, and a drop of 0.19% and 0.21% in the AUC and GAUC metrics on public dataset, respectively. Similarly, after removing HFA, model performance drops by 0.16% in AUC and 0.88% in GAUC on industrial dataset. The main reason is that HFA helps prescreen feature importance, which eventually ensures more effective knowledge extraction in later modules.